# A Survey on Error-Bounded Lossy Compression for Scientific Datasets

### v1

@@ Abstract.
Error-bounded lossy compression has been effective in significantly reducing the data storage/transfer burden while preserving the reconstructed data fidelity very well. Many error-bounded lossy compressors have been developed for a wide range of parallel and distributed use cases for years. These lossy compressors are designed with distinct compression models and design principles, such that each of them features particular pros and cons. In this paper we provide a comprehensive survey of emerging error-bounded lossy compression techniques for different use cases each involving big data to process. The key contribution is fourfold. (1) We summarize an insightful taxonomy of lossy compression into 6 classic compression models. (2) We provide a comprehensive survey of 10+ commonly used compression components/modules used in error-bounded lossy compressors. (3) We provide a comprehensive survey of 10+ state-of-the-art error-bounded lossy compressors as well as how they combine the various compression modules in their designs. (4) We provide a comprehensive survey of the lossy compression for 10+ modern scientific applications and use-cases. We believe this survey is useful to multiple communities including scientific applications, high-performance computing, lossy compression, and big data.
@@ 1. Introduction
Today’s scientific exploration and discovery substantially depend on large-scale scientific simulations or advanced instruments, which can easily produce vast amounts of data. Such vast volumes of data need to be transferred at different levels of devices (such as memory, network, and disk I/O) during the simulations or data acquisition for post hoc analysis. Coherent imaging methods, for example, are one of the primary drivers for the upgrades to the light sources (LCLS-II (lcls-ii, 2017), APS-U (Fornek, 2017), NSLS-II, ALS-U), which will generate high-resolution detector images at a very high frequency, producing data streams of 250 GB/s (Cappello et al., 2019) in some settings. These instrument data are expected to be transferred through an internal high-speed network and stored in a dedicated storage device for later analysis. Another typical example is that 16 petabytes of memory is required to store the full quantum state of a 50-qubit system (Wu et al., 2019).
@@
Error-bounded lossy compression [^1] Data compression is also known as data reduction. In this paper we may use these two terms exchangeably. has been effective in reducing the volumes of scientific datasets for different use cases. Basically, as indicated by our previous study (Cappello et al., 2019), the common use cases that have been explored include significantly reducing storage footprint (Zhao et al., 2022) and memory footprint (Wu et al., 2019), avoiding recomputation cost in scientific simulations (Gok et al., 2018), accelerating checkpoint/restart (Tao et al., 2018), accelerating the I/O performance (Liang et al., 2019b), and reducing data stream intensity (Underwood et al., 2023). More emerging use cases will be discussed in Section 5.
@@
Studies have showed that the reconstructed data of lossy compressors are acceptable to users for their post hoc analysis as long as the compression errors can be controlled to a certain extent. In the community, we call such lossy compressors that allows to control the data distortion to a certain low level error-bounded lossy compressors. Since error-bounded compression can potentially reach very high compression ratios (e.g., 10–1000 (Liang et al., 2018b; Zhao et al., 2021) or even higher (Liu et al., 2023c; Li et al., 2023; Ballester-Ripoll et al., 2019)), this technique is arguably a promising solution to resolve the big data issues in the above-mentioned use cases.
@@
In this paper we present a comprehensive survey to provide a thorough understanding of the error-bounded lossy compression techniques, their pros and cons, and how to use error-bounded lossy compression in different parallel and distributed use cases. The following topics are considered.
  • We summarize an insightful taxonomy of lossy compression into 6 compression models.
@@
  • We provide a comprehensive survey of 10+ commonly used compression components/modules (such as various predictors, bit truncation, quantization, wavelet transform, tucker decomposition, autoencoder) used in different lossy compressors.
@@
  • We choose 30+ state-of-the-art error-controlled lossy compressors (i.e., compression pipelines) and analyze 12 of them in terms of how various compression modules are used in their designs. The studied compressors include not only classic general-purpose error-bounded lossy compressors (such as SZ, ZFP) but also many other emerging tailored lossy compressors (such as SPERR, AESZ, FAZ, MDZ) optimized for specific use cases.
@@
The remainder of the paper is organized as follows. We propose a novel compression model taxonomy in Section 2, which summarizes the existing lossy compression methods into six categories. In Section 3 we survey multiple modular techniques commonly used in different lossy compressors. In Section 4 we discuss existing off-the-shelf lossy compressors for scientific datasets and how they are designed/developed based on the aforementioned lossy compression modules/techniques. In Section 5 we discuss a wide range of applications and the parallel and distributed use cases with regard to lossy compression techniques. We discuss related work in Section 6 and conclude the survey in Section 7 with a discussion of future work.
@@ 2. Compression Model Taxonomy
Figure 1 illustrates the six state-of-the-art data compression models we summarized based on many existing error-bounded lossy compressors. Each of the models has its pros and cons with respect to the time/space complexity and reconstruction quality. In the figure, for example, the models on the left generally tend to have lower time complexity yet lower data reconstruction quality than do the models on the right. [^2] Note that the performance and quality rule demonstrated in Figure 1 holds in general cases. The real performance/quality also depends on the specific design and implementation. In practice, each reduction model is a fundamental technique that can be combined with other models or techniques to generate a specific data compressor.
Figure 1. Error-Controlled Lossy Compression Model Taxonomy
@@
1. Decimation-/filtering-based compression: Decimation can be split into two categories: spatial decimation and temporal decimation. The former often adopts a sampling method during the data compression and then recovers the missing data by applying an interpolation over the sampled data during the data reconstruction. The latter samples the temporal snapshots every $K$ time steps during the simulation or data acquisition and reconstructs the missing snapshots using an interpolation method. Pros: extremely high data compression performance. Cons: potentially very expensive data reconstruction since it needs to recover missing data points by numerical methods such as interpolation. Specific examples will be given in later sections.
@@
3. Transformation-based compression: Data transform (such as wavelet and cosine transform) has been widely used in the data compression community (Lindstrom, 2014; Li et al., 2019; Li, 2018; Li et al., 2023) because it can effectively convert the original data domain to another domain (generally called the coefficient domain). The transformed domain is generally much easier to compress because the coefficient data are mostly close to zero and their values often exhibit regular spatial characteristics (e.g., large values are gathered in a core of the space). Pros: may lead to fairly high rate distortion (i.e., high ratio with high quality); high performance due to matrix multiplication (e.g., on GPU). Cons: not easy to control the error bound; relatively fixed transform methods.
@@
6. Deep-learning-based compression: Deep learning techniques have been used to improve the data compression ratio. In particular, autoencoder (AE) (Goodfellow et al., 2016) and variational autoencoder (VAE) (Kingma and Welling, 2013) are two classic data reconstruction techniques. An autoencoder is a kind of artificial neural network for learning efficient data codings in an unsupervised manner. The original aim of an autoencoder is to learn a representation (encoding) for a set of data, typically for dimensionality reduction. Thus it can be leveraged to reduce data size. By comparison, VAE is a generative model (similar to a generative adversarial network (Goodfellow et al., 2014)), in which a variational approach is used for latent representation learning, resulting in a specific estimator for the training algorithm. VAE has many variate versions (Kingma and Welling, 2013; Kumar et al., 2017; Higgins et al., 2017; Zhao et al., 2017; Tolstikhin et al., 2017; Kolouri et al., 2018; Chen et al., 2018), which are studied in our project. Pros: a fast-emerging technique with a promising opportunity to get a very high compression ratio. Cons: inferior data reconstruction quality; very expensive training; relatively expensive encoding and decoding.
@@ 3. Modular Lossy Compression Techniques
In this section we describe the key lossy compression modules or techniques that are often used in many modern state-of-the-art error-bounded lossy compressors. Each technique listed here generally cannot be treated as a complete compressor but just a critical module or step in a compression pipeline (i.e., a compressor). That is, each technique needs to be combined with one or more other techniques to compose an error-bounded lossy compression pipeline, in order to obtain a high compression ratio with strictly controlled compression errors based on user-defined error bounds.
3.1. Pointwise Data Prediction (PDP)
Data prediction is a critical technique in the prediction-based error-bounded compression model, such as FPZIP and the SZ-series compressors including SZ1.4, SZ2, SZ3, and QoZ, as well as many domain-specific compressors (MDZ, CliZ, etc.). Generally in the whole compression pipeline, the data prediction step is the first or second step, followed by computing the difference between the predicted value and the original value, which would lead to a set of close-to-zero values. These close-to-zero values could be compressed more easily/effectively than the original data values.
@@ 3.1. Pointwise Data Prediction (PDP)
Figure 2. RRS Policy with Six Prediction Methods
Table 1. Predictors Used in Different Lossy Compressors
@@
Scaled-Pattern Pastri (Gok et al., 2018) Quantum Che. Many
Temporal Smoothness MDZ (Zhao et al., 2022) MD 1
Multi-level MDZ (Zhao et al., 2022) MD Many
Mask-based CliZ (Jian et al., 2024) Climate 2 or 4
3.2. Quantization (QT)
Quantization is a popular technique widely used in today’s lossy compressors. Basically, quantization means a specific procedure/operation in which the data value range will be split into multiple consecutive intervals (i.e., quantization bins) each with a unique bin number, and each data value would be checked in which bin/interval it is located so that it can be represented by the corresponding quantization bin number. After the quantization step, each quantization bin would contain a certain number of data values, forming a histogram, which would often be encoded by a certain coding algorithm (such as Huffman encoding) to get a high compression ratio. The data to be quantized could be the original data values (Laboratory, 2023; Zhang et al., 2023) or the difference of the predicted value and original data value (Tao et al., 2017c; Liang et al., 2018b). Various types of quantization methods are summarized in Table 2 and are detailed in the following text.
Table 2. Quantizations Used in Different Lossy Compressors
@@ 3.2. Quantization (QT)
Vector-quantization MDZ MD Matching multilevel pattern
@@
  • Vector quantization. Similar to log-scale quantization, vector quantization adopts variable-length quantization bins, where the quantization bin size depends on a certain clustering (e.g., K-means) technique applied on the dataset. This can improve the data approximation accuracy when using the centroid to represent all the data values contained by the corresponding bins. Typical examples that use the vector quantization method include MDZ (Zhao et al., 2022) and NUMARCK (Chen et al., 2014).
@@
3.3. Orthogonal/Wavelet Transform (OWT/DWT)
@@
3.4. Pointwise Domain Transform (PDT)
Pointwise domain transform here refers to a (pre)processing step that performs an operation on each data point in order to meet a specific error bound requirement. A typical example of PDT is using logarithmic domain transform to implement the pointwise relative error bound (Zhao et al., 2020a). Specifically, Liang et al. (Liang et al., 2018a) proved that enforcing a pointwise relative error bound $e_{r}$ on the original data $d$ is equivalent to enforcing an absolute error bound $\log(1+e_{r})$ on the logarithmic data $\log|d|$. Thus, compression with pointwise relative error bound can be implemented as traditional lossy compression with absolute error bound after performing a logarithmic transform on the original data. This is a generic approach that can be applied to any compressor with an absolute error bound. However, it will introduce certain overhead due to the expensive logarithmic operations during compression and exponential operations during decompression.
3.5. Bit-Plane Coding (BPC)
Bit-plane coding is commonly used in many lossy compressors with different compression models. BPC can be applied either in the original data domain or in the transformed coefficient domain. The fundamental idea about BPC is that the scientific data are always stored in a specific bit-plane representation (e.g., IEEE 754 floating-point or integer), such that each bit in the presentation affects the data value with different levels. Taking 32-bit floating-point data as an example (as illustrated in Figure 3), the alteration of leading bits (high-end) will change the data value more significantly than the alteration of ending bits (low-end). The reason is that the leading part contains the sign, exponent, and significant mantissa bits. Thus, ignoring a certain number of insignificant bit planes for a group of data values is often used in different lossy compression algorithms. In general, the loss introduced into the data by the bit truncation method is determined by the data values: the larger the data value, the larger the compression error. We explain the reason by using a floating-point value as an example. For simplicity and without loss of generality, we give the analysis based on the floating-point value in decimal format instead of binary format actually used by the data representation on machines. For the two numbers 12.34 and 123.4, their representations are 1.234$\times$$10^{1}$ and 1.234$\times$$10^{2}$. When a digit is removed (e.g., removing 4), the errors introduced into the two numbers would be 0.04 and 0.4, respectively, which depends on the original data values. Performing BPC after aligning the exponent of data to the same scale is a typical variation, which enforces an absolute error bound that is irrelevant to the data value.
Figure 3. Example of Bit Truncation to Remove 6 Insignificant Bits for 5 Consecutive Data Values

### v2

@@ Abstract.
Error-bounded lossy compression has been effective in significantly reducing the data storage/transfer burden while preserving the reconstructed data fidelity very well. Many error-bounded lossy compressors have been developed for a wide range of parallel and distributed use cases for years. They are designed with distinct compression models and principles, such that each of them features particular pros and cons. In this paper we provide a comprehensive survey of emerging error-bounded lossy compression techniques. The key contribution is fourfold. (1) We summarize a novel taxonomy of lossy compression into 6 classic models. (2) We provide a comprehensive survey of 10 commonly used compression components/modules. (3) We summarized pros and cons of 46 state-of-the-art lossy compressors and present how state-of-the-art compressors are designed based on different compression techniques. (4) We discuss how customized compressors are designed for specific scientific applications and use-cases. We believe this survey is useful to multiple communities including scientific applications, high-performance computing, lossy compression, and big data.
@@ 1. Introduction
Today’s scientific exploration and discovery substantially depend on large-scale scientific simulations or advanced instruments, which can easily produce vast amounts of data. Such vast volumes of data need to be transferred at different levels of devices (such as memory, network, and disk I/O) during the simulations or data acquisition for post hoc analysis. Coherent imaging methods, for example, are one of the primary drivers for the upgrades to the light sources (LCLS-II (lcl, [n. d.]), APS-U (Fornek, 2017), NSLS-II, ALS-U), which will generate high-resolution detector images at a very high frequency, producing data streams of 250 GB/s (Cappello et al., 2019) in some settings. Another typical example is that 16 petabytes of memory is required to store the full quantum state of a 50-qubit system (Wu and et al., 2019).
@@
Lossy compression [^1] Data compression is also known as data reduction. In this paper we may use these two terms exchangeably. has been effective in reducing the volumes of scientific data for different use cases. The scientific data to compress are generally referred to as the datasets used/generated by scientific applications or instruments, which are often stored in floating-point or integer values. Basically, as indicated by our previous study (Cappello et al., 2019), the common use cases that have been explored include significantly reducing storage footprint (Zhao et al., 2022) and memory footprint (Wu and et al., 2019), avoiding recomputation cost in scientific simulations (Gok et al., 2018), accelerating checkpoint/restart (Tao et al., 2018), accelerating the I/O performance (Liang et al., 2019b), and reducing data stream intensity (Underwood et al., 2023b). More emerging use cases will be discussed in Section 6.
@@
Studies have showed that the reconstructed data of lossy compressors are acceptable to users for their post hoc analysis as long as the compression errors can be controlled to a certain extent. Such lossy compressors that allows to control the data distortion are often called error-bounded lossy compressors. Since error-bounded compression can potentially reach very high compression ratios (e.g., 10–1000 (Liang et al., 2018b; Zhao et al., 2021) or even higher (Liu et al., 2023c; Li et al., 2023; Ballester-Ripoll et al., 2019)), this technique is arguably a promising solution to resolve the big data issues for scientific applications.
@@
In this paper we present a comprehensive survey to provide a thorough understanding of the error-controlled lossy compression techniques and how they are used in different parallel and distributed use cases. The following topics are considered.
  • We summarize a novel taxonomy of lossy compression into 6 compression models.
@@
  • We provide a comprehensive survey of 10 commonly used compression components/modules (such as various predictors, bit truncation, quantization, wavelet transform, tucker decomposition, autoencoder) used in different lossy compressors.
@@
  • We summarize 46 state-of-the-art lossy compressors (i.e., compression pipelines) and use the representative compressors to describe how compression modules are used in the compression design. The studied compressors include not only classic general-purpose error-bounded lossy compressors (such as SZ, ZFP) but also many emerging tailored lossy compressors (such as SPERR, AESZ, FAZ, MDZ) optimized for specific use cases.
@@
The remainder of the paper is organized as follows. We discuss related work in Section 2. We propose a compression model taxonomy in Section 3. In Section 4 we survey modular techniques commonly used in lossy compressors. In Section 5 we discuss off-the-shelf lossy compressors for scientific datasets and how they are developed based on the aforementioned lossy compression modules/techniques. In Section 6 we discuss a wide range of applications and the parallel and distributed use cases. We conclude the survey in Section 7 with a discussion of future work.
2. Related Work
Compression survey across domains and data types.
  • Jayasankar et al. (Jayasankar et al., 2021) contributed a comprehensive survey to summarize the data compression techniques in terms of different coding schemes (such as entropy coding and dictionary coding) amd across various data types (such as text compression, image compression, audio compression, and video compression). The survey also involved various use cases, including compression for wireless sensor networks, medical imaging, database compression, HEP data compression, and wind turbine data compression.
  • Son et al. (Son et al., 2014) provided a survey about data compression for scientific domains that were generally produced by HPC applications. This survey involves both lossless compression techniques (such as FPC (Burtscher and Ratanaworabhan, 2009), ISOBAR (Schendel and et al., 2012), and PRIMACY (Shah and et al., 2012)) and four lossy compressors (such as ISABELA (Lakshminarasimhan et al., 2011) and fpzip (Lindstrom et al., 2017)).
  • There are many survey studies about image compression (to name a few (Jamil et al., 2023; Hussain et al., 2018)). These surveys focus on the compression for the image data. The major techniques include various transforms such as discrete cosine trans- form (DCT) and discrete wavelet transforms (DWT) wavelet transform, learning-driven image compression (Jamil et al., 2023) such as autoencoder (AE), variational autoencoder (VAE), and convolutional neural network (CNN).
The above surveys have limited information about modern lossy compression for scientific datasets. For example, they missed state-of-the-art (SOTA) error-bounded lossy compressors such as SZ (Di and Cappello, 2016; Tao et al., 2017c), ZFP (Lindstrom, 2014), MGARD (Ainsworth et al., 2018), SPERR (Li et al., 2023), and TTHRESH (Ballester-Ripoll et al., 2019). Our survey provides an in-depth survey of lossy compression for scientific datasets, which features error-control or data distortion control in terms of scientific analysis. We also develop a compression model taxonomy, and present compression techniques and how lossy compressors are used in applications.
Compression Survey for Specific Domains or Use Cases
  • Time series data: Chiarot et al. (Chiarot and Silvestri, 2023) provided a survey about the principal time series compression techniques, proposing a taxonomy to classify them considering their overall approach and their characteristics. The authors also discussed the performance of the selected algorithms by comparing the experimental results that were provided in the original articles.
Table 1. Brief comparison of survey works on data compression.
Survey Focus Limitation
Jayasankar et al. General compression Lacks error-bounded compressors
Son et al. HPC scientific data Lacks SOTA techniques
Image Surveys Image compression Limited to image compression
Mummadisetty et al. Climate data Focuses on lossless methods
Hilal et al. Seismic data Limited generalizability
Quach et al. Point cloud compression Limited generalizability
Our Survey Scientific datasets Comprehensive, fills prior gaps
In comparison, our survey has the following unique features: (1) a comprehensive lossy compression model taxonomy (Section 3), (2) discussing modular compression techniques (Section 4), (3) summary of 46 general-purpose error-controlled lossy compressors (Section 5), and (4) comprehensively discussing compressor customization for specific applications (Section 6). We present a comparison in Table 1.
@@ 3. Compression Model Taxonomy
Figure 1. Scientific Lossy Compression Model Taxonomy
We developed a taxonomy derived from 46 general-purpose scientific lossy compressors (listed in Table 6), categorizing them into six state-of-the-art lossy compression models, as depicted in Figure 1. These six compression models are outlined using a unified compression pipeline, shown in Figure 2. The steps enclosed in blue-bordered boxes represent the most critical components of each model, which serve as the basis for naming the corresponding compression model.
Figure 2. Compression Pipeline with Various Models: each highlighted box represents the key step of the corresponding model. All the compression techniques shown here will be detailed in Section 4.
In Figure 1, each model has its pros and cons with respect to the time/space complexity and reconstruction quality. The models on the left generally tend to have lower time complexity yet lower data reconstruction quality than do the models on the right. [^2] Note that the performance and quality rule demonstrated in Figure 1 holds in general cases. The real performance/quality also depends on the specific design and implementation. In practice, each reduction model is a fundamental technique that can be combined with other models or techniques to generate a specific data compressor, which will be detailed in Section 5.
@@
1. Decimation-/filtering-based compression: Decimation can be split into two categories: spatial decimation and temporal decimation. The former often adopts a sampling method during the data compression and then recovers the missing data by applying an interpolation over the sampled data during the data reconstruction. The latter samples the temporal snapshots every $K$ time steps during the simulation or data acquisition and reconstructs the missing snapshots using an interpolation method. Pros: extremely high data compression performance. Cons: potentially very expensive data reconstruction since it needs to recover missing data points by numerical methods such as interpolation. Specific examples will be given in Section Section 4.7.
@@
3. Transformation-based compression: Data transform (such as wavelet and cosine transform) has been widely used in the data compression community (Lindstrom, 2014; Li et al., 2019; Li, 2018; Li et al., 2023) because it can effectively decorrelate data by converting the original data domain to another so-called coefficient domain. The transformed domain is easier to compress as most coefficients are near zero and exhibit spatial regularity, with large values clustered at a corner of the space. For instance, the Haar wavelet in SSEM (Sasaki et al., 2015) calculates deltas between adjacent data points across dimensions, often yielding many near-zero values due to the dataset’s high smoothness. Pros: may lead to fairly high rate distortion (i.e., high ratio with high quality); high performance due to matrix multiplication (e.g., on GPU). Cons: not easy to control the error bound; relatively fixed transform methods.
@@
6. Deep-learning-based compression: Deep learning techniques have been used to improve the data compression ratio. In particular, autoencoder (AE) (Goodfellow et al., 2016) and variational autoencoder (VAE) (Kingma and Welling, 2013; Kumar et al., 2017; Higgins et al., 2017; Zhao et al., 2017; Tolstikhin et al., 2017; Kolouri et al., 2018; Chen et al., 2018) are two classic data reconstruction techniques. An autoencoder is a kind of artificial neural network for learning efficient data codings in an unsupervised manner. The original aim of an autoencoder is to learn a representation (encoding) for a set of data, typically for dimensionality reduction. Pros: a fast-emerging technique with a promising opportunity to get a very high compression ratio. Cons: inferior data reconstruction quality; very expensive training; relatively expensive encoding and decoding.
@@ 4. Modular Lossy Compression Techniques
Figure 3. Usage Distribution of Compression Techniques in 46 Compressors (e.g., 25.7% compressors in Table 6 used LE).
In this section we describe the key lossy compression modules/techniques often used in many modern state-of-the-art lossy compressors. Each technique listed here serves as just a module or step in a compression pipeline (i.e., a compressor). That is, each technique often needs to be combined with one or more other techniques to compose an error-bounded lossy compression pipeline in order to obtain a high compression ratio, forming a many-to-many relationship between compression-pipeline and compression-technique. Figure 3 presents the usage distribution of compression techniques in the 46 lossy compressors we collected.
4.1 Pointwise Data Prediction (PDP)
Data prediction is a critical technique in the prediction-based error-bounded compression model, such as FPZIP and the SZ-series compressors including SZ1.4, SZ2, SZ3, and QoZ, as well as many domain-specific compressors (MDZ (Zhao et al., 2022), CliZ (Jian et al., 2024), etc.). Generally in the whole compression pipeline, the data prediction step is the first or second step, followed by computing the difference between the predicted value and the original value, which would lead to a set of close-to-zero values. These close-to-zero values could be compressed more easily/effectively than the original data values.
@@ 4.1 Pointwise Data Prediction (PDP)
Figure 4. RRS Policy with Six Prediction Methods. SZ1, SZ2, SZ3, SPERR are general-purpose compressors (see Table 6); Pastri (Gok et al., 2018) and MDZ (Zhao et al., 2022) are customized compressors for chemistry applications (Sec. 6.1 & 6.2).
@@
Table 2. Predictors Used in Lossy Compressors: the last column indicates the number of neighbor data values used in that predictor. The details about the predictors can be found in the corresponding compressor papers.
@@
Scaled-Pattern Pastri (Gok et al., 2018) (Section 6.2) Quantum Che. Many
Temporal Smoothness MDZ (Zhao et al., 2022) (Section 6.1) MD 1
Multi-level MDZ (Zhao et al., 2022) (Section 6.1) MD Many
Mask-based CliZ (Jian et al., 2024) (Section 6.4) Climate 2 or 4
4.2 Quantization (QT)
Table 3. Quantizations Used in Lossy Compressors: different quantization methods feature specific approximation effect as shown in the last column.
@@ 4.2 Quantization (QT)
Vector-quantization MDZ/NUMARCK MD Matching multilevel pattern
@@
Quantization is a popular technique widely used in today’s lossy compressors. Basically, quantization means a specific procedure/operation in which the data value range will be split into multiple consecutive intervals (i.e., quantization bins) each with a unique bin number, and each data value would be checked in which bin/interval it is located so that it can be represented by the corresponding quantization bin number. After the quantization step, each quantization bin would contain a certain number of data values, forming a histogram, which would often be encoded by a certain coding algorithm (such as Huffman encoding) to get a high compression ratio. The data to be quantized could be the original data values (Huang et al., 2023a; Zhang et al., 2023) or the difference of predicted value and original data (Tao et al., 2017c; Liang et al., 2018b). We summarize various types of quantization in Table 3.
@@
  • Vector quantization. Similar to log-scale quantization, vector quantization adopts variable-length quantization bins, where the quantization bin size depends on a certain clustering (e.g., K-means) technique applied on the dataset. This can improve the data approximation accuracy when using the centroid to represent all the data values contained by the corresponding bins. Typical examples that use the vector quantization include MDZ (Zhao et al., 2022) and NUMARCK (Chen et al., 2014).
@@
4.3 Orthogonal/Wavelet Transform (OWT/DWT)
@@
4.4 Domain Transform (DT) Domain transform here refers to a (pre)processing step that performs an operation on the dataset in order to meet a specific error bound requirement. A typical example of DT is using logarithmic domain transform to implement the pointwise relative error bound (Zhao et al., 2020a). Specifically, Liang et al. (Liang et al., 2018a) proved that enforcing a pointwise relative error bound $e_{r}$ on the original data $d$ is equivalent to enforcing an absolute error bound $\log(1+e_{r})$ on the logarithmic data $\log|d|$. Thus, compression with pointwise relative error bound can be implemented as traditional lossy compression with absolute error bound after performing a logarithmic transform on the original data. This is a generic approach that can be applied to any compressor with an absolute error bound. However, it will introduce certain overhead due to the expensive logarithmic operations during compression and exponential operations during decompression. Another good example about DT is zMesh (Luo et al., 2021), which reorder the adaptive mesh refinement (AMR) dataset by an optimized space-filling curve. This can significantly improve the smoothness of the data, which thus improve the compression ratio substantially in turn.
4.5 Bit-Plane Coding (BPC)
Bit-plane coding is commonly used in many lossy compressors with different compression models. BPC can be applied either in the original data domain or in the transformed coefficient domain. The fundamental idea about BPC is that the scientific data are always stored in a specific bit-plane representation (e.g., IEEE 754 floating-point or integer), such that each bit in the presentation affects the data value with different levels.
Taking 32-bit floating-point data as an example, the alteration of leading bits (high-end) will change the data value more significantly than the alteration of ending bits (low-end). The reason is that the leading part contains the sign, exponent, and significant mantissa bits. Thus, ignoring a certain number of insignificant bit planes for a group of data values is often used in different lossy compression algorithms. In general, the loss introduced into the data by the bit truncation method is determined by the data values: the larger the data value, the larger the compression error. We explain the reason by using a floating-point value as an example. For simplicity and without loss of generality, we give the analysis based on the floating-point value in decimal format instead of binary format actually used by the data representation on machines. For the two numbers 12.34 and 123.4, their representations are 1.234$\times$$10^{1}$ and 1.234$\times$$10^{2}$. When a digit is removed (e.g., removing 4), the errors introduced into the two numbers would be 0.04 and 0.4, respectively, which depends on the original data values. Performing BPC after aligning the exponent of data to the same scale is a typical variation, which enforces an absolute error bound that is irrelevant to the data value.